OpenStudy (anonymous):

JamesJ: would i be able to do the same with this problem??? integral 0-pi/4 sec^2 t dt

OpenStudy (across):

I like how you're thinking, rmalik2. :)

OpenStudy (anonymous):

what you mean..?

OpenStudy (anonymous):

INT sec^t dt = tan t + c between limits 0 and pi/4 forget c and calculate tan pi/4 - tan 0

OpenStudy (across):

I assume you're asking JamesJ if you can do this:\[\int x^n dx=\frac{x^{n+1}}{n+1}\]with this\[\int_{0}^{\pi}4\cdot\sec^2(t) dt.\]

OpenStudy (jamesj):

So the 'trick' here is knowing how to integrate sec^2(x)

OpenStudy (anonymous):

yea i know the identity of sec^2t = 1-tan^2t

OpenStudy (jamesj):

which is a bit subtle. What is the derivative of f(x) = tan x ?

OpenStudy (anonymous):

yes lol

OpenStudy (jamesj):

btw, your identity isn't quite right. 1 + tan^2 x = sec^2 x

OpenStudy (anonymous):

o sorry 1 + tan^2 x = sex ^2x

OpenStudy (jamesj):

Now: What is the derivative of f(x) = tan x ?

OpenStudy (jamesj):

calculate using the quotient rule: tan x = sin x / cos x

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

f(x)= -log cosx?

OpenStudy (jamesj):

no, you integrated. What is the derivative of tan x = sin x / cos x ?

OpenStudy (jamesj):

Use the quotient rule

OpenStudy (anonymous):

im not sure

OpenStudy (jamesj):

Write u = sin x and v = cos x Then what is the derivative of u/v ?

OpenStudy (jamesj):

Do you know the quotient rule?

OpenStudy (anonymous):

no i dont

OpenStudy (jamesj):

Do you know the product rule?

OpenStudy (jamesj):

the product rule for differentiating, that is.

OpenStudy (anonymous):

i think so would it be the chain rule??

OpenStudy (jamesj):

No, that's another thing.

OpenStudy (jamesj):

The product rule says if f is function that can be written as f(x) = u(x).v(x) for some other functions u and v. Then if u and v can be differentiated, then the derivative of f is given by f' = u'.v + u.v'

OpenStudy (anonymous):

o well then im not sure

OpenStudy (jamesj):

The quotient rule says if we can write f = u/v then the derivative of f is \[f' = \frac{v.u' - u.v'}{v^2}\]

OpenStudy (jamesj):

Do these look familiar now?

OpenStudy (anonymous):

oooooooo yea now i get that

OpenStudy (jamesj):

Ok. Now write tan x = sin x / cos x. What is its derivative?

OpenStudy (anonymous):

but im not sure

OpenStudy (jamesj):

Let u = sin x and v = cos x. Now apply the quotient rule that I just wrote down.

OpenStudy (jamesj):

I know this feels like a long way from your problem, but trust me, we're on the right road and once we turn the corner it will be clear.

OpenStudy (jamesj):

Got it yet?

OpenStudy (jamesj):

If u = sin x and v = cos x What are u' and v' ?

OpenStudy (anonymous):

u' is -cosx v' sinx

OpenStudy (jamesj):

Right. So now apply the quotient rule

OpenStudy (anonymous):

to where?

OpenStudy (jamesj):

tan x = sin x / cos x = u/v And if f = u/v, then \[f' = \frac{v.u' - u.v'}{v^2}\]

OpenStudy (jamesj):

I.e., calculate the derivative of tan x.

OpenStudy (anonymous):

o so it becomes sinx * -cosx- cos x *sin x/ sin^2x

OpenStudy (jamesj):

No, you've got everything backwards. Look again at what u and v are.

OpenStudy (anonymous):

well i did say u'=-cosx and v'=sinx...so (cos x)(-cos x)-(sinx)(sinx)/cos^2x?

OpenStudy (jamesj):

Oh, there' the problem. If u = sin x, u' = cos x If v = cos x, v' = -sin x

OpenStudy (anonymous):

i thought its -cos x

OpenStudy (jamesj):

Hence for f(x) = tan x \[f' = \frac{\cos x . \cos x - \sin x.(-\sin x)}{\cos^2 x} = \frac{\cos^2 x + \sin^2 x}{\cos^2 x} = \frac{1}{\cos^2 x}\]

OpenStudy (jamesj):

i.e., f'(x) = sec^2 x.

OpenStudy (jamesj):

Hence applying this result to your original problem and using the fundamental theorem of calculus, the indefinite integral of sec^2 x is tan x.

OpenStudy (anonymous):

yes but why go through all this when i know that it would be tan pi/4 -tan(0)=1?

OpenStudy (jamesj):

The question is HOW do you know that the integral of sec^2 x is tan x?

OpenStudy (jamesj):

...because the problems won't always be this easy and you need to see how this works so when a sophisticated variation comes along, you can figure it out.

OpenStudy (anonymous):

well sinx+ cosx=1 and when i divid sinx i would get 1+tanx=1/secx right?

OpenStudy (jamesj):

What?!

OpenStudy (jamesj):

sin^2 x + cos^2 x = 1

OpenStudy (anonymous):

yes i meant that i just didnt put the ^2 on

OpenStudy (jamesj):

and you divide by cos^2 x to get tan^2 x + 1 = sec^2 x

OpenStudy (jamesj):

But this last identity has nothing to do with your problem.

OpenStudy (anonymous):

ok